# Deep Generative Data Assimilation in Multimodal Setting

### v1

@@ 1 Introduction
Data assimilation (DA) is crucial for numerous scientific disciplines that require accurate modeling of chaotic systems. It is particularly important across various areas of the geosciences, including fluid dynamics simulations, oceanography, and atmospheric science [8]. At its core, the state estimation problem of DA seeks to calibrate the background trajectory of states, $x^{b}_{1:T}=(x^{b}_{1},\cdots,x^{b}_{T})\in\mathbb{R}^{T\times N_{x}}$, given available observations, $y$, which are often noisy, sparse, and/or multimodal. The calibrated state after assimilation, denoted as $x^{a}_{1:T}$, is referred to as the analysis (see Figure 1).
@@ 2.1 Multimodality in Unified Latent Space
Fortunately, with deep learning, we are able to circumvent this problem by learning a sufficiently expressive function that is able to project both the states and observations into a unified, reduced-order $\mathcal{L}$-dimensional latent space. In this work, we define $f_{\theta}:\mathbb{R}^{(N_{x}+M_{1}+\cdots+M_{K})}\rightarrow\mathbb{R}^{\mathcal{L}}$ as an encoder that takes as input the concatenation of background state and each observation modality along unique channels for a single timestep. The decoder $g_{\phi}:\mathbb{R}^{\mathcal{L}}\rightarrow\mathbb{R}^{N_{x}}$ then maps the latent code back into the pixel-space. For notational simplicity, and for the rest of the paper, we adopt $x$ and $y$ to denote the latent representations of the unconditional and conditioning inputs, respectively. Thus, the inverse problem, aligning with Equation 1 is formally considered in the latent space.
@@ Acknowledgments
The authors would like to acknowledge funding from the NSF LEAP Science and Technology Center award #2019625, Department of Energy grant.

### v2

@@ 1 Introduction
Data assimilation (DA) is crucial for numerous scientific disciplines that require accurate modeling of chaotic systems. It is particularly important across various areas of the geosciences, including fluid dynamics simulations, oceanography, and atmospheric science [8]. At its core, the state estimation problem of DA seeks to calibrate background trajectory of states, $x^{b}_{1:T}=(x^{b}_{1},\cdots,x^{b}_{T})\in\mathbb{R}^{T\times N_{x}}$, given available observations, $y$, which are often noisy, sparse, and/or multimodal. The calibrated state after assimilation, denoted as $x^{a}_{1:T}$, is referred to as the analysis (see Figure 1).
@@ 2.1 Multimodality in Unified Latent Space
Fortunately, with deep learning, we are able to circumvent this problem by learning a sufficiently expressive function that is able to project both the states and observations into a unified, reduced-order $\mathcal{L}$-dimensional latent space. In this work, we define $f_{\theta}:\mathbb{R}^{(N_{x}+M_{1}+\cdots+M_{K})}\rightarrow\mathbb{R}^{\mathcal{L}}$ as an encoder that takes as input the concatenation of background state and each observation modality along unique channels for a single timestep. The decoder $g_{\phi}:\mathbb{R}^{\mathcal{L}}\rightarrow\mathbb{R}^{N_{x}}$ then maps the latent code back into the pixel-space.
In the following section, we review previous work on score-based diffusion model and its novel application to DA introduced by [47]. In our extension, we define variables $x$ and $y$ to represent the latent representations of the unconditional and conditioning inputs, respectively, including in-situ and ex-situ observations to align with Equation 1.
@@ Acknowledgments
The authors would like to acknowledge funding from the National Science Foundation Science and Technology Center, Learning the Earth with Artificial intelligence and Physics, LEAP (Grant number 2019625), Department of Energy grant.